Solve for x and y
6x-3y=9
5x+2y=-15

Answer by susie(3)

to use the elimination method:
first you must get rid of one variable (x or y) using multiplication
2(6x-3y=9)
3(5x+2y=-15)

12x-6y=18 (then you add the 2 equations
15x+6y=-45 which cancels out the y variable)
+ -----------
27x = -27 (then divide both sides by 27 to isolate the x)
so x = -1
(then plug x into either equation to find y)

12(-1)-6y=18
-12-6y=18 (add 12 to both sides to isolate y)

you end up with -6y=30 (then divide both sides by -6)

so y=-5 x=-1 (you can plug both into an equation to check your answer)